OpenStudy (anonymous):

Julie kicks a moose skin ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the moose skin ball hits the ground again? 0.6 sec 0.8 sec 1.6 secs 2.8 secs please help...

OpenStudy (cwrw238):

use the equations of motion with constant accleeration

OpenStudy (cwrw238):

s = ut + 0.5gt^2

OpenStudy (anonymous):

what are the equations if you dont mind:)

OpenStudy (cwrw238):

s = displacement = 0 g = -32 ft/s62 and u = 25 ft/s

OpenStudy (cwrw238):

0 = 25t - 16t^2 solve for t

OpenStudy (anonymous):

cwrw238 you are so wrong that newton cam back to life and died again

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

so what would the answer be

OpenStudy (anonymous):

Michelle i would like to help you but i think that the data is incomplete...

OpenStudy (anonymous):

but that is how the question is written out what am suppose to do? i could go back and reread the lesson

OpenStudy (anonymous):

yes you should you dont even know what equations of motion are

OpenStudy (anonymous):

you see this is my third time taking this class and i really don't want to fail again.

OpenStudy (anonymous):

Ahaha michelle take it easy!

OpenStudy (anonymous):

i think I will, look over the lesson again,but what do you guys think would be the best way to approach it.

OpenStudy (anonymous):

with love

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

The problem that i am having is the fact that they didnt mention at which point in height the balls stops and starts falling back

OpenStudy (anonymous):

yeah, i guess, i could always email my teacher about it because there are alway's glitches like that happening in the tests.

jimthompson5910 (jim_thompson5910):

cwrw238 already said it solve 0 = 25t - 16t^2 for t to get your answer (it will be the root that is not 0)

OpenStudy (anonymous):

um ok, i will go back and do it that way.

OpenStudy (anonymous):

thanks for the advice everyone, you all have been great so you have a nice day

OpenStudy (anonymous):

YOU HAVE TO USE SEQENCES EVERTHING ELSE IS WRONG!!!!!!!!!!!!!!

OpenStudy (anonymous):

Good luck!

OpenStudy (cwrw238):

16t^2 - 25 = 0 t(16t - 25) = 0 16t = 25 f = 25/16 = 1.6